# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 3.1 Ecology and epidemiology
Maciel et al. introduced nonlocal terms in the competitive terms in the Lotka-Volterra competition model (25). They studied the effect of the range of nonlocal interaction on each species’ influence ranges maciel2021 . In this case, we have two scenarios: additive and non-additive influence ranges. In the case of non-additive influence ranges, individuals affect competitors of the same or other species in a spatial range that depends only on the influence range of the individual itself. An application of this could be the allelopathic interactions between plants, in which the extension of the roots or branches determines the species’ influence range bais2003 ; bertin2003 ; cheng2015 . Another application could be the different types of antagonistic interactions established in microbial communities and governed by the release of chemicals into the environment granato2019 ; czaran2002 ; nadell2016 . On the other hand, for additive influence ranges, the ranges of the nonlocal interactions are a combination of the influence ranges of the species engaged in the interaction. This signifies that the plants compete for resources through their roots or home-ranging animals competing for forage over a finite region, such as central place foragers houston1985 ; olsson2008 .
@@
The mixed hyperbolic-parabolic nonlocal model (26) has two important key features colombo2014 . First, the prey moves in all directions due to the presence of Laplacian term, and predators move to regions with higher prey density. Second, the range of nonlocal interaction of nonlocal terms defines how far predators can “feel” the presence of prey, and they move accordingly.
@@
In the classical SIR model, it is assumed that the epidemic is in a homogeneous population kermack1927 . But, generally, a disease in an epidemic spreads heterogeneously among individuals. So, a nonlocal version (13) of the SIR model is an advancement model to investigate the propagation of the epidemic kendall1965 . Researchers have been studying diffusion and nonlocal dispersal with delayed transmission in the SIR model wang2015 ; tian2017 ; liu2019 ; zhou2020 . In kendall1965 , Kendall proved the existence of a travelling wave solution of this nonlocal model with a minimum speed. Another type of nonlocal epidemic model has been studied in recent days kang2021 ; kang2020 ; kang2021b , namely age-structured susceptible-infectious-susceptible (SIS). This is mainly a nonlocal spatial diffusion model, slightly different from earlier considerations. Suppose, at time $t\geq 0$ and at $x\in\Omega\subset\mathbb{R}^{N}$, $S(t,a,x)$ and $I(t,a,x)$ are the densities of susceptible and infective individuals of age $a\geq 0$. We further assume that $a^{+}$ is the maximum of the individuals and is finite. Now, for $t>0$, $a\in(0,a^{+}]$, and $x\in\Omega$, we consider an age-structured nonlocal model as kang2021b :
@@
These models consider nonlocal interactions between individuals, accounting for long-range disease transmission and spatial interactions that traditional local epidemiological models do not capture. These are essential in scenarios where geographic distances influence the spread of infections. Researchers have been studying the nonlocal dispersal epidemic models with and without age structures kang2021 ; kang2020 ; kang2021b ; xu2020 ; yang2019 ; yang2013 . Most of these studies have been focused on the existence of travelling wave solutions to capture the transmission of the disease. In addition, researchers have studied the memory effect on disease propagation with the help of FODE to reveal the inherent memory of the disease propagation based on the size of the initial infected population saeedian2017 .
Different types of models can be employed to study the spread of infectious diseases, and they may include elements such as population dynamics, epidemiological parameters, and the effects of interventions like vaccination. As discussed in the previous section, the transmission of the COVID-19 virus is nonlocal because people travelling between different areas have the potential to bring the virus to new places or contribute to its spread within and between groups. One of the most important strategies for halting the COVID-19 outbreak has been vaccination shi2023 . Effective vaccinations limit the overall spread of the virus among the community by reducing the severity of the illness and promoting herd immunity. These transmissions can be simulated and predicted using mathematical models, which often incorporate differential equations. These models can account for variables, including travel patterns, vaccination coverage, recovery rates, and infection rates. Researchers use computational models to investigate various scenarios and provide guidance for public health policies. For instance, compartmental models (e.g., nonlocal SIR models discussed earlier), agent-based models, and spatial models are often used for this purpose. Policymakers can use these models to inform choices on vaccination distribution, travel bans, and other ways to stop the virus from spreading.
@@ 3.2 Biology, medicine and health sciences
Nonlocal interactions in cell biology are an exciting field of study because they provide a deeper understanding of cell behaviours and tissue dynamics in a variety of healthy and pathological circumstances. It gives more insight into the cooperation and exchange of information among cells in complex biological systems as cellular imaging and computational tools advance. Nonlocal interaction refers to the long-range signalling and communication between cells that extend beyond their immediate neighbours, and it captures various biological phenomena in cell biology, such as tissue morphogenesis chen2020 and wound healing webb2022 . In addition, cancer cells interact with the extracellular matrix and surrounding tissues in tumour growth and invasion over extended distances bitsouni2018 . Therefore, these models have been employed to describe tumor growth patterns and the influence of micro-environmental factors. Furthermore, nonlocal interaction plays a significant role in coordinating the movements of cells within the migrating group, enabling efficient collective migration.
@@
The spatial spreading patterns of invasive cell populations have been modelled by the Fisher-KPP equation and its extensions and applied in vivo malignant spreading swanson2008 ; swanson2003 ; perez2018 and in vitro cell biology experiments vo2015 ; warne2019 ; tremel2009 ; sengers2007 ; cai2007 . The most common disease, ‘cancer’, is a collection of diseases described by uncontrolled growth of cells and tumor development that invades the tissue of origin and distant organs hillen2013 . Researchers have suggested that intrinsic and environmental bottlenecks challenge early tumour growth and progression folkman1991 ; d2005 ; teng2008 ; almog2009 ; hanahan2011 . In the initial stage, a small subset of cancer cells called ‘cancer stem cells’ intrinsically populates the tumour. The agent-based nonlocal model (6) describes the time evolution of cancer stem cells and tumour cells hillen2013 . This modelling approach widely applies to solid, avascular tumours in any organ.
@@
The nonlocal model (9) can be applied to various aspects of cell migration and growth chen2020 . For example, this model could account for the tumour cell heterogeneity behaviour influenced by the tumour microenvironment’s composition. The memory effect is one of the key factors in disease propagation, and in this case, time-fractional RD equations (e.g., the nonlocal model (10)) are used in the mathematical model to incorporate such an evolution process.
@@
In addition, anomalous diffusion processes, characterized by fractional diffusion equations, have been observed in biological transport systems, such as the movement of molecules within cells or through porous media. These models also have been applied to study the movement of molecules, proteins, and lipids in biological membranes. The anomalous transport accounts for the nonlocal interactions and memory effects of diffusing species within the membrane, e.g., tumor growth and invasion solis2019 . These models further capture the anomalous transport of cancer cells and the interactions between tumor cells and their microenvironment, providing insights into tumor growth patterns and invasion dynamics. In addition, these models account for the nonlocal and memory effects of drug diffusion within the polymer, providing a more accurate representation of drug release kinetics yin2011 . Furthermore, the nonlocal models with fractional derivatives have been applied to describe anomalous transport of molecules in biological tissues, such as fractional Fick’s law for describing fractional diffusion of solutes in biological membranes.
@@
Fractional models are used to describe the movement of neurotransmitters and other molecules within neurons, accounting for the nonlocal interactions and subdiffusive behaviour observed in neuronal dynamics magin2010 . Furthermore, these models are used to describe the memory effects and long-range interactions in immune cell recruitment and activation during an immune response.
@@
Biology conservation laws represent fundamental mass or energy conservation principles in biological systems. They ensure that the total amount of mass or energy within a specific domain remains constant over time, which is essential for the accurate modelling of biological processes. For instance, conservation laws in biology include mass conservation in biological transport sterner2011 , conservation of energy in metabolic processes pang2014 ; idumah2023 , and conservation of species or population densities in ecological systems santini2022 .
@@
Following the basic principles of aggregation and dispersal, we can divide the velocity $\mathbf{p}$ into two groups: individuals move towards and away from each other mogilner1999 ; edelstein1998 ; topaz2004 ; bertozzi2009 . Suppose $\mathbf{p}_{a}$ and $\mathbf{p}_{d}$ are the attractive and dispersive velocities of the populations at time $t\geq 0$, and hence the resultant velocity $\mathbf{p}=\mathbf{p}_{a}+\mathbf{p}_{d}$. The species’ aggregation may occur due to some sensing mechanism (e.g., smell or sight), and we can assume a spatial average of nearby populations with some weights for this sensing, i.e., nonlocal sensing. Taking all these into consideration, the attractive nonlocal velocity as topaz2006 :
@@
$V$ is the species-specific attractive speed, $l$ is the characteristic range of the nonlocal sensing, and $\alpha$ is the characteristic density. On the other hand, dispersal arises from the anti-crowding mechanism, which operates over a shorter scale mogilner1999 ; breder1954 , and we can consider this as spatially local. The direction of dispersal is generally opposite to the direction of population gradients, and the simplest model in describing these effects as
@@
Following topaz2006 , after rescaling this model, we obtain:
@@
This dimensionless model (43) conserves the zeroth and first moments of $u$ but does not conserve higher moments topaz2006 . The zeroth moment shows that the total mass is conserved, while the first moment implies that the centre of mass is conserved.
Now, we consider a nonlocal chemotaxis model for cell migration, which follows the mass conservation law. Suppose $p=p(t,\mathbf{x},v,\widehat{\mathbf{v}})$ is the density distribution of a cell population at a mesoscopic level at time $t>0$ and position $\mathbf{x}\in\Omega\subset\mathbb{R}^{d}$ conte2022 . Here $v$ is the statistical distribution of the speed $v\in[0,V]$ with the maximal speed $V$, and $\widehat{\mathbf{v}}\in\mathbb{S}^{d-1}$ is the polarization directions. Then, the transport equation for the cell distribution at the mesoscopic level is given by conte2022 :
@@
where $\mathcal{F}[p](t,\mathbf{x},v,\widehat{\mathbf{v}})$ is the scattering of the microscopic velocity in direction and speed. The run and tumble are the typical microscopic dynamics of a cell berg1983 ; block1983 . These are modelled by the velocity jump process, characterized by a turning frequency $\mu$ and transition probability $T$ conte2022 ; stroock1974 . Assuming the nonlocal transition probability loy2020 and its independency on the prep-orientation velocity stroock1974 ; othmer1988 ; hillen2006 , the turning operator becomes conte2022 :
@@
In the two-dimensional bounded domain, authors in conte2022 studied the model (44) with mass conservation. By incorporating nonlocal interactions and conservation laws, mathematical models in biology can better capture the complexities of biological systems and provide deeper insights into the behaviour of living organisms. These modelling approaches are particularly relevant in studying systems with spatial dependencies, memory effects, and long-range interactions, which are prevalent in many biological processes.
@@
Nonlocal interactions in cell biology refer to interactions between cellular components that extend beyond immediate neighbours and may involve long-range signalling or communication. These interactions play a crucial role in cellular processes and behaviours, such as cell migration, pattern formation, and collective cell responses. Cellular adhesion is one of the fundamental forms of cell-cell interactions, and its continuum modelling remains mathematically challenging buttenschon2018 . The adhesion model (20) is the first continuum model, verified experimentally in Steinberg’s adhesion-driven cell-sorting experiments armstrong2006 . This model has been extended and applied to many areas, e.g., cell-cell and cell-extracellular matrix (ECM) adhesion in studying the invasion of cancer cells into the ECM gerisch2008 ; sherratt2009 ; painter2010 ; andasari2011 ; chaplain2011 ; domschke2014 . This type of nonlocal advection-reaction-diffusion model can capture the cell-cell interactions to account for chemotactic movements due to overcrowding or the action of the stellates’ processes on hepatocytes armstrong2006 ; sherratt2009 ; mogilner1995 ; green2010 . The coexistence of chemotaxis occurs in vivo in disease propagation, e.g., wound healing or cancer progression bromberek2002 ; conte2022 . Fibroblasts migrate efficiently along collagen or fibronectin fibers in connective tissues during wound healing provenzano2009 . In the case of tumor spread, cancer cells follow the aligned fibers at the tumor-stroma interface. The nonlocal model (44) for cell distributions can capture cell migration at the mesoscopic level.
@@
Different types of self-organization biological aggregations exist in nature, e.g., insect swarms, ungulate herds, fish schools, and bacterial colonies ben2000 ; camazine2020 . These aggregations arise from different social phenomena without the influence of external stimuli such as attraction, group cohesion, repulsion, and collision avoidance mogilner1999 . Sometimes, biological aggregation forms distinct groups with a sharp edge, called clumping parrish1997 ; parrish1999 . This clumping behaviour can be observed analytically levine2000 and numerically topaz2004 for two or higher-dimensional models but not for one-dimensional models built on realistic assumptions about social interactions on swarming behaviours. On the other hand, the clumping behaviour can arise in a simple nonlocal biological aggregation model (43) for one or higher-dimensional models. The model (43) preserves integro-differential conservation law with two movement terms: nonlinear degenerate diffusion arising from anti-crowding behaviour, and the other describes attractive nonlocal social interactions topaz2006 . Further research into nonlocal interactions in cell biology will help to a greater understanding of cellular processes and their significance to development, illness, and tissue engineering as experimental tools and computational methodologies evolve.
@@
Energy-harvesting electronics and gadgets have gained popularity in biomedical applications, having the potential to power medical equipment and sensors utilizing ambient energy sources. For example, electrical energy may be produced from mechanical vibrations or motions of bodily motion. In addition, the temperature gradients within the body or the surrounding environment can be used to generate electricity. Energy-harvesting technologies provide implanted medical devices with a long-term power source, such as triboelectric or piezoelectric generators, which can eliminate the need for battery replacement krishnaswamy2020c ; wang2017 ; buroni2020 ; krishnaswamy2019 ; krishnaswamy2019a ; krishnaswamy2019b ; krishnaswamy2020 ; wang2018 ; krishnaswamy2020a ; krishnaswamy2020b ; krishnaswamy2021 ; krishnaswamy2023 ; buroni2024 . Implantable devices driven by harvested energy provide precise and regulated medication delivery, increasing the efficacy of medicinal therapies zou2021 . The development of self-powered sensors is essential in remote patient monitoring systems and ensures continuous data collection without the need for frequent maintenance yang2013a .

### v2

@@ 3.1 Ecology and epidemiology
Maciel et al. introduced nonlocal terms in the competitive terms in the Lotka-Volterra competition model (24). They studied the effect of the range of nonlocal interaction on each species’ influence ranges [209]. In this case, we have two scenarios: additive and non-additive influence ranges. In the case of non-additive influence ranges, individuals affect competitors of the same or other species in a spatial range that depends only on the influence range of the individual itself. An application of this could be the allelopathic interactions between plants, in which the extension of the roots or branches determines the species’ influence range [211, 362, 363]. Another application could be the different types of antagonistic interactions established in microbial communities and governed by the release of chemicals into the environment [212, 364, 365]. On the other hand, for additive influence ranges, the ranges of the nonlocal interactions are a combination of the influence ranges of the species engaged in the interaction. This signifies that the plants compete for resources through their roots or home-ranging animals competing for forage over a finite region, such as central place foragers [366, 367].
@@
The mixed hyperbolic-parabolic nonlocal model (25) has two important key features [213]. First, the prey moves in all directions due to the presence of Laplacian term, and predators move to regions with higher prey density. Second, the range of nonlocal interaction of nonlocal terms defines how far predators can “feel” the presence of prey, and they move accordingly.
@@
In the classical SIR model, it is assumed that the epidemic is in a homogeneous population [15]. But, generally, a disease in an epidemic spreads heterogeneously among individuals. So, a nonlocal version (12) of the SIR model is an advancement model to investigate the propagation of the epidemic [57, 368]. Researchers have studied diffusion and nonlocal dispersal with delayed transmission in the SIR model [369, 370, 371, 372]. In [57], Kendall proved the existence of a travelling wave solution of this nonlocal model with a minimum speed. Another type of nonlocal epidemic model has been studied in recent days [58, 373, 374], namely age-structured susceptible-infectious-susceptible (SIS). This is mainly a nonlocal spatial diffusion model, slightly different from earlier considerations. Suppose, at time $t\geq 0$ and at $x\in\Omega\subset\mathbb{R}^{N}$, $S(t,a,x)$ and $I(t,a,x)$ are the densities of susceptible and infective individuals of age $a\geq 0$. We further assume that $a^{+}$ is the maximum of the individuals and is finite. Now, for $t>0$, $a\in(0,a^{+}]$, and $x\in\Omega$, we consider an age-structured nonlocal model as [374]:
@@
These models consider nonlocal interactions between individuals, accounting for long-range disease transmission and spatial interactions that traditional local epidemiological models do not capture. These are essential in scenarios where geographic distances influence the spread of infections. Researchers have been studying the nonlocal dispersal epidemic models with and without age structures [58, 373, 374, 375, 376, 377]. Most of these studies have been focused on the existence of travelling wave solutions to capture the transmission of the disease. In addition, researchers have studied the memory effect on disease propagation with the help of FODE to reveal the inherent memory of the disease propagation based on the size of the initial infected population [24].
Different types of models can be employed to study the spread of infectious diseases, and they may include elements such as population dynamics, epidemiological parameters, and the effects of interventions like vaccination. As discussed in the previous section, the transmission of the COVID-19 virus is nonlocal because people travelling between different areas have the potential to bring the virus to new places or contribute to its spread within and between groups. One of the most important strategies for halting the COVID-19 outbreak has been vaccination [140]. Effective vaccinations limit the overall spread of the virus among the community by reducing the severity of the illness and promoting herd immunity. These transmissions can be simulated and predicted using mathematical models, which often incorporate differential equations. These models can account for variables, including travel patterns, vaccination coverage, recovery rates, and infection rates. Researchers use computational models to investigate various scenarios and provide guidance for public health policies. For instance, compartmental models (e.g., nonlocal SIR models discussed earlier), agent-based models, and spatial models are often used for this purpose. Policymakers can use these models to inform choices on vaccination distribution, travel bans, and other ways to stop the virus from spreading.
@@ 3.2 Biology, medicine and health sciences
Nonlocal interactions in cell biology are an exciting field of study because they provide a deeper understanding of cell behaviours and tissue dynamics in a variety of healthy and pathological circumstances. It gives more insight into the cooperation and exchange of information among cells in complex biological systems as cellular imaging and computational tools advance. Nonlocal interaction refers to the long-range signalling and communication between cells that extend beyond their immediate neighbours, and it captures various biological phenomena in cell biology, such as tissue morphogenesis [52] and wound healing [378]. In addition, cancer cells interact with the extracellular matrix and surrounding tissues in tumour growth and invasion over extended distances [379]. Therefore, these models have been employed to describe tumor growth patterns and the influence of micro-environmental factors. Furthermore, nonlocal interaction plays a significant role in coordinating the movements of cells within the migrating group, enabling efficient collective migration.
@@
The spatial spreading patterns of invasive cell populations have been modelled by the Fisher-KPP equation and its extensions and applied in vivo malignant spreading [380, 381, 382] and in vitro cell biology experiments [383, 384, 385, 386, 387]. The most common disease, ‘cancer’, is a collection of diseases described by uncontrolled growth of cells and tumor development that invades the tissue of origin and distant organs [110]. Researchers have suggested that intrinsic and environmental bottlenecks challenge early tumour growth and progression [388, 389, 390, 391, 392]. In the initial stage, a small subset of cancer cells called ‘cancer stem cells’ intrinsically populates the tumour. The agent-based nonlocal model (6) describes the time evolution of cancer stem cells and tumour cells [110]. This modelling approach widely applies to solid, avascular tumours in any organ.
@@
The nonlocal model (8) can be applied to various aspects of cell migration and growth [52]. For example, this model could account for the tumour cell heterogeneity behaviour influenced by the tumour microenvironment’s composition. The memory effect is one of the key factors in disease propagation, and in this case, time-fractional RD equations (e.g., the nonlocal model (9)) are used in the mathematical model to incorporate such an evolution process.
@@
In addition, anomalous diffusion processes, characterized by fractional diffusion equations, have been observed in biological transport systems, such as moving molecules within cells or through porous media [393]. These models also have been applied to study the movement of molecules, proteins, and lipids in biological membranes. The anomalous transport accounts for the nonlocal interactions and memory effects of diffusing species within the membrane, e.g., tumor growth and invasion [394]. These models further capture the anomalous transport of cancer cells and the interactions between tumor cells and their microenvironment, providing insights into tumor growth patterns and invasion dynamics. In addition, these models account for the nonlocal and memory effects of drug diffusion within the polymer, providing a more accurate representation of drug release kinetics [395]. Furthermore, the nonlocal models with fractional derivatives have been applied to describe anomalous transport of molecules in biological tissues, such as fractional Fick’s law for describing fractional diffusion of solutes in biological membranes.
@@
Fractional models are used to describe the movement of neurotransmitters and other molecules within neurons, accounting for the nonlocal interactions and subdiffusive behaviour observed in neuronal dynamics [396]. Furthermore, these models are used to describe the memory effects and long-range interactions in immune cell recruitment and activation during an immune response.
@@
Biology conservation laws represent fundamental mass or energy conservation principles in biological systems. They ensure that the total amount of mass or energy within a specific domain remains constant over time, which is essential for the accurate modelling of biological processes. For instance, conservation laws in biology include mass conservation in biological transport [397], conservation of energy in metabolic processes [398, 399], scalar conservation law in nonlocal traffic flow model [400], and conservation of species or population densities in ecological systems [401].
@@
Following the basic principles of aggregation and dispersal, we can divide the velocity $\mathbf{p}$ into two groups: individuals move towards and away from each other [293, 402, 403, 404]. Suppose $\mathbf{p}_{a}$ and $\mathbf{p}_{d}$ are the attractive and dispersive velocities of the populations at time $t\geq 0$, and hence the resultant velocity $\mathbf{p}=\mathbf{p}_{a}+\mathbf{p}_{d}$. The species’ aggregation may occur due to different sensing mechanisms (e.g., smell or sight), and we can assume a spatial average of nearby populations with some weights for this sensing, i.e., nonlocal sensing. Taking all these into consideration, the attractive nonlocal velocity as [302]:
@@
$V$ is the species-specific attractive speed, $l$ is the characteristic range of mechanismscal sensing, and $\alpha$ is the characteristic density. On the other hand, dispersal arises from the anti-crowding mechanism, which operates over a shorter scale [293, 405], and we can consider this as spatially local. The direction of dispersal is generally opposite to the direction of population gradients, and the simplest model in describing these effects as
@@
Following [302], after rescaling this model, we obtain:
@@
This dimensionless model (51) conserves the zeroth and first moments of $u$ but does not conserve higher moments [302]. The zeroth moment shows that the total mass is conserved, while the first moment implies that the centre of mass is conserved.
Now, we consider a nonlocal chemotaxis model for cell migration, which follows the mass conservation law. Suppose $p=p(t,\mathbf{x},v,\widehat{\mathbf{v}})$ is the density distribution of a cell population at a mesoscopic level at time $t>0$ and position $\mathbf{x}\in\Omega\subset\mathbb{R}^{d}$ [406]. Here $v$ is the statistical distribution of the speed $v\in[0,V]$ with the maximal speed $V$, and $\widehat{\mathbf{v}}\in\mathbb{S}^{d-1}$ is the polarization directions. Then, the transport equation for the cell distribution at the mesoscopic level is given by [406]:
@@
where $\mathcal{F}[p](t,\mathbf{x},v,\widehat{\mathbf{v}})$ is the scattering of the microscopic velocity in direction and speed. The run and tumble are the typical microscopic dynamics of a cell [407, 408]. These are modelled by the velocity jump process, characterized by a turning frequency $\mu$ and transition probability $T$ [406, 409]. Assuming the nonlocal transition probability [410] and its independency on the prep-orientation velocity [409, 411, 412], the turning operator becomes [406]:
@@
In the two-dimensional bounded domain, authors in [406] studied the model (52) with mass conservation. By incorporating nonlocal interactions and conservation laws, mathematical models in biology can better capture the complexities of biological systems and provide deeper insights into the behaviour of living organisms. These modelling approaches are particularly relevant in studying systems with spatial dependencies, memory effects, and long-range interactions, which are prevalent in many biological processes.
@@
Nonlocal interactions in cell biology refer to interactions between cellular components that extend beyond immediate neighbours and may involve long-range signalling or communication. These interactions play a crucial role in cellular processes and behaviours, such as cell migration, pattern formation, and collective cell responses. Cellular adhesion is one of the fundamental forms of cell-cell interactions, and its continuum modelling remains mathematically challenging [153, 413]. The adhesion model (18) is the first continuum model, verified experimentally in Steinberg’s adhesion-driven cell-sorting experiments [154]. This model has been extended and applied to many areas, e.g., cell-cell and cell-extracellular matrix (ECM) adhesion in studying the invasion of cancer cells into the ECM [414, 155, 415, 416, 417, 418]. This type of nonlocal advection-reaction-diffusion model can capture the cell-cell interactions to account for chemotactic movements due to overcrowding or the action of the stellates’ processes on hepatocytes [154, 155, 419, 420]. The coexistence of chemotaxis occurs in vivo in disease propagation, e.g., wound healing or cancer progression [421, 406]. Fibroblasts migrate efficiently along collagen or fibronectin fibers in connective tissues during wound healing [422]. Cancer cells follow the aligned fibers at the tumor-stroma interface in the case of tumour spread. The nonlocal model (52) for cell distributions can capture cell migration at the mesoscopic level. The nonlocal single equation model resulting from the full chemotaxis system is often obtained by simplifying the complex system of PDEs that explain chemotaxis, which is the movement of cells or organisms in response to chemical gradients [423]. This reduction frequently occurs in research on chemotaxis systems’ pattern formation, aggregation, and blow-up processes.
@@
Different types of self-organization biological aggregations exist in nature, e.g., insect swarms, ungulate herds, fish schools, and bacterial colonies [424, 425, 313, 298, 299, 56]. These aggregations arise from different social phenomena without the influence of external stimuli such as attraction, group cohesion, repulsion, and collision avoidance [293]. Sometimes, biological aggregation forms distinct groups with a sharp edge, called clumping [426, 427]. This clumping behaviour can be observed analytically [49] and numerically [403] for two or higher-dimensional models but not for one-dimensional models built on realistic assumptions about social interactions on swarming behaviours. On the other hand, the clumping behaviour can arise in a simple nonlocal biological aggregation model (51) for one or higher-dimensional models. The model (51) preserves integro-differential conservation law with two movement terms: nonlinear degenerate diffusion arising from anti-crowding behaviour, and the other describes attractive nonlocal social interactions [302]. Further research into nonlocal interactions in cell biology will help to a greater understanding of cellular processes and their significance to development, illness, and tissue engineering as experimental tools and computational methodologies evolve.
@@
Energy-harvesting electronics and gadgets have gained popularity in biomedical applications, having the potential to power medical equipment and sensors utilizing ambient energy sources. For example, electrical energy may be produced from mechanical vibrations or bodily motions. In addition, the temperature gradients within the body or the surrounding environment can be used to generate electricity. Energy-harvesting technologies provide implanted medical devices with a long-term power source, such as triboelectric or piezoelectric generators, which can eliminate the need for battery replacement [282, 428, 429, 430, 431, 432, 433, 434, 435, 436, 437, 438, 439]. Implantable devices driven by harvested energy provide precise and regulated medication delivery, increasing the efficacy of medicinal therapies [440]. The development of self-powered sensors is essential in remote patient monitoring systems and ensures continuous data collection without the need for frequent maintenance [441].